OpenStudy (anonymous):

A car of mass 1200kg is travelling at 12ms-1. When the brakes are applied the car comes uniformly to rest in 6.0seconds. (a) Calculate the kinetic energy lost by the car.

OpenStudy (anonymous):

Ek= 1/2 m v^2

OpenStudy (anonymous):

Ok. So 1/2*1200*?. How would i calculate v ?

OpenStudy (anonymous):

its speed of 12 ms-1

OpenStudy (anonymous):

Is that V^2 ?

OpenStudy (anonymous):

Ek= 1/2 * 1200 kg * (12 ms-1)^2 = 86400 J = 86,4 kJ

OpenStudy (anonymous):

Ok. Great